OpenStudy (anonymous):

Solve 5s + 40 = 55 using number sense. Please help! Thanks! :)

OpenStudy (mathmale):

"Number sense" ?? Start by subtracting 40 from both sides of the given equation.

OpenStudy (anonymous):

Simplifying 5s + 40 = 55 Reorder the terms: 40 + 5s = 55 Solving 40 + 5s = 55 Solving for variable 's'. Move all terms containing s to the left, all other terms to the right. Add '-40' to each side of the equation. 40 + -40 + 5s = 55 + -40 Combine like terms: 40 + -40 = 0 0 + 5s = 55 + -40 5s = 55 + -40 Combine like terms: 55 + -40 = 15 5s = 15 Divide each side by '5'. s = 3 Simplifying s = 3

OpenStudy (anonymous):

I know, I don't get "number sense" either. Thanks! :) I'll make sure to become a fan! :)